a book club promises to send 8 books for $1, if you join the club. after you recieve 8 books, you may select more books at a rate of $19.99 per book. if you spend a total of $80.96, how many extra books did you purchase
Answer by askmemath(368)

Total cost = $80.96
Membership fee = $1
So total spent on books = 80.96 - 1 = 79.96
Cost of each book = $19.99
Total number of books bought = 79.96/19.99 = 4
So you purchased 4 more books